# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

### v1

@@ 6. Conclusion
This study comprehensively explores how Indian users integrate PHI wearables for health management. Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of PHI in India. By doing so, we emphasize the influence of sociocultural factors, usability challenges, and trust issues in their consistent/long-term use of PHI for health management. While PHI and specifically wearables hold significant potential to shift personal health management from a reactive to a proactive model, their integration is hindered by cultural norms, limited health literacy, and skepticism around data accuracy. To address these challenges, we propose a dual focus on improving health literacy and building trust through transparent communication, medical validation, and user-centered design.
Two essential contributions of this research are identifying critical areas where designers and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants. The critical areas include expanding PHI functionalities to align with family-based caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information. As shown by the evaluated prototype, the HCI community, technology developers, and policymakers must work together to standardize core functionalities and integrate PHI into formal healthcare systems. By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive health management, supporting both individual users and familial care structures.
This work offers a framework for future research and industry efforts, reinforcing the transformative potential of PHI within diverse cultural contexts. By addressing technological and sociocultural barriers, PHI can aid in preventive healthcare, fostering long-term engagement and broader adoption across India and beyond.
@@ Acknowledgements.
We would like to thank the student Harsh Pala for his involvement in the development of Figma prototypes used for evaluating an interactive interface of the sharing preferences identified by the card sorting activity.References
@@
- Jones et al. (2021) Simon L. Jones, William Hue, Ryan M. Kelly, Rosemarie Barnett, Violet Henderson, and Raj Sengupta. 2021. Determinants of Longitudinal Adherence in Smartphone-Based Self-Tracking for Chronic Health Conditions: Evidence from Axial Spondyloarthritis. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 5, 1 (March 2021), 16:1–16:24. https://doi.org/10.1145/3448093
@@
- Lee et al. (2024a) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee. 2024a. FamilyScope: Visualizing Affective Aspects of Family Social Interactions Using Passive Sensor Data. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 57:1–57:27. https://doi.org/10.1145/3637334
- Lee et al. (2024b) Jarrett G.W. Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe. 2024b. Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2 (May 2024), 51:1–51:21. https://doi.org/10.1145/3659585
@@
- Poole (2013) Erika S. Poole. 2013. HCI and Mobile Health Interventions. Translational Behavioral Medicine 3, 4 (Dec. 2013), 402–405. https://doi.org/10.1007/s13142-013-0214-3

### v2

@@ 6. Conclusion
This study comprehensively explores how Indian users integrate PHI wearables for proactive care. Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of this approach in India. By doing so, we emphasize the influence of sociocultural factors (collective care practices), usability challenges (fragmented ecosystems), and trust issues (lack of provenance) in their consistent/long-term use of PHI for proactive care. Two essential contributions of this research are identifying critical areas where designers, developers, and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants. The critical areas include expanding PHI functionalities to align with collective caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information. As demonstrated by the evaluated prototype and the recommendations, the HCI community, technology developers, and policymakers must collaborate to standardize core functionalities and integrate PHI into formal healthcare systems. By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive care, supporting both individual users and collective care structures.
@@ Acknowledgements.
We would like to thank student Harsh Pala for his involvement in the development of Figma prototypes used to evaluate the interactive interface of the sharing preferences identified through the card sorting activity.References
@@
- Aneja et al. (2025) Urvashi Aneja, Aarushi Gupta, and Aditya Vashistha. 2025. Beyond Semantics: Examining Gender Bias in LLMs Deployed within Low-resource Contexts in India. In Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25). Association for Computing Machinery, New York, NY, USA, 2784–2795. https://doi.org/10.1145/3715275.3732180
@@
- Bathelt et al. (2022) Franziska Bathelt, Ines Reinecke, Brita Sedlmayr, Sepp Höhne, Christian Gierschner, and Martin Sedlmayr. 2022. Visualization of medical wearable-data using SMART-on-FHIR–A concept for an interdisciplinary complex practical course. In Digital Professionalism in Health and Care: Developing the Workforce, Building the Future. IOS Press, Amsterdam, Netherlands, 61–65.
@@
- Binda et al. (2018) Jomara Binda, Chien Wen Yuan, Natalie Cope, Hyehyun Park, Eun Kyoung Choe, and John M. Carroll. 2018. Supporting Effective Sharing of Health Information among Intergenerational Family Members. In Proceedings of the 12th EAI International Conference on Pervasive Computing Technologies for Healthcare (New York, NY, USA) (PervasiveHealth ’18). Association for Computing Machinery, New York, NY, USA, 148–157. https://doi.org/10.1145/3240925.3240936
@@
- Cha et al. (2025) Yoon Jeong Cha, Jiongyu Chen, Yasemin Gunal, Qiying Zhu, Mark W Newman, and Sun Young Park. 2025. Collaborative Health-Tracking Technologies for Children and Parents: A Review of Current Studies and Directions for Future Research. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 1038, 13 pages. https://doi.org/10.1145/3706598.3713596
@@
- Chang et al. (2024) Zhaoxin Chang, Fusang Zhang, Xujun Ma, Pei Wang, Weiyan Chen, Duo Zhang, Badii Jouaber, and Daqing Zhang. 2024. MmECare: Enabling Fine-grained Vital Sign Monitoring for Emergency Care with Handheld MmWave Radars. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 4, Article 207 (Nov. 2024), 24 pages. https://doi.org/10.1145/3699766
- Chen et al. (2024) Wenqiang Chen, Jiaxuan Cheng, Leyao Wang, Wei Zhao, and Wojciech Matusik. 2024. Sensor2Text: Enabling Natural Language Interactions for Daily Activity Tracking Using Wearable Sensors. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 4, Article 192 (Nov. 2024), 26 pages. https://doi.org/10.1145/3699747
@@
- Cho et al. (2021) Sylvia Cho, Chunhua Weng, Michael G Kahn, and Karthik Natarajan. 2021. Identifying Data Quality Dimensions for Person-Generated Wearable Device Data: Multi-Method Study. JMIR Mhealth Uhealth 9, 12 (23 Dec 2021), e31618. https://doi.org/10.2196/31618
- Chopra et al. (2025) Shaan Chopra, Katherine Juarez, James Fogarty, and Sean A. Munson. 2025. Engagements with Generative AI and Personal Health Informatics: Opportunities for Planning, Tracking, Reflecting, and Acting around Personal Health Data. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 75 (Sept. 2025), 33 pages. https://doi.org/10.1145/3749503
- Choube et al. (2025) Akshat Choube, Ha Le, Jiachen Li, Kaixin Ji, Vedant Das Swain, and Varun Mishra. 2025. GLOSS: Group of LLMs for Open-ended Sensemaking of Passive Sensing Data for Health and Wellbeing. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 76 (Sept. 2025), 32 pages. https://doi.org/10.1145/3749474
- Chung et al. (2017) Chia-Fang Chung, Nanna Gorm, Irina A. Shklovski, and Sean Munson. 2017. Finding the Right Fit: Understanding Health Tracking in Workplace Wellness Programs. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 4875–4886. https://doi.org/10.1145/3025453.3025510
@@
- Epstein et al. (2020) Daniel A. Epstein, Clara Caldeira, Mayara Costa Figueiredo, Xi Lu, Lucas M. Silva, Lucretia Williams, Jong Ho Lee, Qingyang Li, Simran Ahuja, Qiuer Chen, Payam Dowlatyari, Craig Hilby, Sazeda Sultana, Elizabeth V. Eikey, and Yunan Chen. 2020. Mapping and Taking Stock of the Personal Informatics Literature. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 4, 4, Article 126 (Dec. 2020), 38 pages. https://doi.org/10.1145/3432231
@@
- Feustel et al. (2018) Clayton Feustel, Shyamak Aggarwal, Bongshin Lee, and Lauren Wilcox. 2018. People Like Me: Designing for Reflection on Aggregate Cohort Data in Personal Informatics Systems. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 2, 3, Article 107 (Sept. 2018), 21 pages. https://doi.org/10.1145/3264917
@@
- Gupta et al. (2022) Meghna Gupta, Devansh Mehta, Anandita Punj, and Indrani Medhi Thies. 2022. Sophistication with Limitation: Understanding Smartphone Usage by Emergent Users in India. In Proceedings of the 5th ACM SIGCAS/SIGCHI Conference on Computing and Sustainable Societies (Seattle, WA, USA) (COMPASS ’22). Association for Computing Machinery, New York, NY, USA, 386–400. https://doi.org/10.1145/3530190.3534824
@@
- Harrison et al. (2015) Daniel Harrison, Paul Marshall, Nadia Bianchi-Berthouze, and Jon Bird. 2015. Activity tracking: barriers, workarounds and customisation. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Osaka, Japan) (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 617–621. https://doi.org/10.1145/2750858.2805832
@@
- Intille (2004) Stephen S Intille. 2004. Ubiquitous computing technology for just-in-time motivation of behavior change. In MEDINFO 2004. IOS Press, Amsterdam, Netherlands, 1434–1437.
- Jiang et al. (2023) Zhihan Jiang, Lin Lin, Xinchen Zhang, Jianduo Luan, Running Zhao, Longbiao Chen, James Lam, Ka-Man Yip, Hung-Kwan So, Wilfred H.S. Wong, Patrick Ip, and Edith C.H. Ngai. 2023. A Data-Driven Context-Aware Health Inference System for Children during School Closures. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 1, Article 18 (March 2023), 26 pages. https://doi.org/10.1145/3580800
@@
- Jones et al. (2021) Simon L. Jones, William Hue, Ryan M. Kelly, Rosemarie Barnett, Violet Henderson, and Raj Sengupta. 2021. Determinants of Longitudinal Adherence in Smartphone-Based Self-Tracking for Chronic Health Conditions: Evidence from Axial Spondyloarthritis. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 5, 1, Article 16 (March 2021), 24 pages. https://doi.org/10.1145/3448093
@@
- Kaur and Singh (2025) Jasmeet Kaur and Pushpendra Singh. 2025. Exploring the Potential of Peer Support Group for Family Planning Needs of Women in Resource-constrained Settings in India. Proc. ACM Hum.-Comput. Interact. 9, 2, Article CSCW080 (May 2025), 25 pages. https://doi.org/10.1145/3710978
@@
- Ketabdar and Polzehl (2009) Hamed Ketabdar and Tim Polzehl. 2009. Fall and emergency detection with mobile phones. In Proceedings of the 11th International ACM SIGACCESS Conference on Computers and Accessibility (Pittsburgh, Pennsylvania, USA) (Assets ’09). Association for Computing Machinery, New York, NY, USA, 241–242. https://doi.org/10.1145/1639642.1639695
@@
- Kirchner et al. (2021) Susanne Kirchner, Jessica Schroeder, James Fogarty, and Sean A. Munson. 2021. “They don’t always think about that”: Translational Needs in the Design of Personal Health Informatics Applications. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 210, 16 pages. https://doi.org/10.1145/3411764.3445587
@@
- Ko et al. (2010) Jeonggil Ko, Jong Hyun Lim, Yin Chen, Rvăzvan Musvaloiu-E, Andreas Terzis, Gerald M. Masson, Tia Gao, Walt Destler, Leo Selavo, and Richard P. Dutton. 2010. MEDiSN: Medical emergency detection in sensor networks. ACM Trans. Embed. Comput. Syst. 10, 1, Article 11 (Aug. 2010), 29 pages. https://doi.org/10.1145/1814539.1814550
@@
- Le et al. (2024) Ha Le, Rithika Lakshminarayanan, Jixin Li, Varun Mishra, and Stephen Intille. 2024. Collecting Self-reported Physical Activity and Posture Data Using Audio-based Ecological Momentary Assessment. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 3, Article 111 (Sept. 2024), 35 pages. https://doi.org/10.1145/3678584
- Lee et al. (2024a) Hyunsoo Lee, Yugyeong Jung, Hei Yiu Law, Seolyeong Bae, and Uichin Lee. 2024a. PriviAware: Exploring Data Visualization and Dynamic Privacy Control Support for Data Collection in Mobile Sensing Research. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 787, 17 pages. https://doi.org/10.1145/3613904.3642815
- Lee et al. (2024b) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee. 2024b. FamilyScope: Visualizing Affective Aspects of Family Social Interactions Using Passive Sensor Data. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 57:1–57:27. https://doi.org/10.1145/3637334
- Lee et al. (2024c) Jarrett G.W. Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe. 2024c. Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2, Article 51 (May 2024), 21 pages. https://doi.org/10.1145/3659585
- Lee et al. (2024d) Jarrett G.W. Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe. 2024d. Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2 (May 2024), 51:1–51:21. https://doi.org/10.1145/3659585
@@
- Li et al. (2010) Ian Li, Anind Dey, and Jodi Forlizzi. 2010. A stage-based model of personal informatics systems. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 557–566. https://doi.org/10.1145/1753326.1753409
- Li et al. (2021) Qingyang Li, Clara Caldeira, Daniel A. Epstein, and Yunan Chen. 2021. Supporting Caring among Intergenerational Family Members through Family Fitness Tracking. In Proceedings of the 14th EAI International Conference on Pervasive Computing Technologies for Healthcare (Atlanta, GA, USA) (PervasiveHealth ’20). Association for Computing Machinery, New York, NY, USA, 1–10. https://doi.org/10.1145/3421937.3422018
@@
- Lin et al. (2025) Georgianna Lin, Minh Ngoc Le, Khai N. Truong, and Alex Mariakakis. 2025. The Cognitive Strategies Behind Multimodal Health Sensemaking: A Menstrual Health Tracking Case Study. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 110 (Sept. 2025), 27 pages. https://doi.org/10.1145/3749482
- Lin et al. (2024) Georgianna Lin, Brenna Li, Jin Yi Li, Chloe Zhao, Khai Truong, and Alexander Mariakakis. 2024. Users’ Perspectives on Multimodal Menstrual Tracking Using Consumer Health Devices. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 3, Article 116 (Sept. 2024), 24 pages. https://doi.org/10.1145/3678575
- Loerakker et al. (2024) Meagan B. Loerakker, Jasmin Niess, Marit Bentvelzen, and Paweł W. Woźniak. 2024. Designing Data Visualisations for Self-Compassion in Personal Informatics. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 4, Article 169 (Jan. 2024), 22 pages. https://doi.org/10.1145/3631448
- Lu et al. (2020) Lin Lu, Jiayao Zhang, Yi Xie, Fei Gao, Song Xu, Xinghuo Wu, and Zhewei Ye. 2020. Wearable Health Devices in Health Care: Narrative Systematic Review. JMIR Mhealth Uhealth 8, 11 (9 Nov 2020), e18907. https://doi.org/10.2196/18907
@@
- Mandl et al. (2020) Kenneth D Mandl, Daniel Gottlieb, Joshua C Mandel, Vladimir Ignatov, Raheel Sayeed, Grahame Grieve, James Jones, Alyssa Ellis, and Adam Culbertson. 2020. Push button population health: the SMART/HL7 FHIR bulk data access application programming interface. NPJ digital medicine 3, 1 (2020), 151.
@@
- Moore et al. (2022) Jimmy Moore, Pascal Goffin, Jason Wiese, and Miriah Meyer. 2022. An Interview Method for Engaging Personal Data. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 5, 4, Article 173 (Dec. 2022), 28 pages. https://doi.org/10.1145/3494964
- Motahar and Wiese (2022) Tamanna Motahar and Jason Wiese. 2022. A Review of Personal Informatics Research for People with Motor Disabilities. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 6, 2, Article 66 (July 2022), 31 pages. https://doi.org/10.1145/3534614
@@
- Niess et al. (2021) Jasmin Niess, Paweł W. Woźniak, Yomna Abdelrahman, Passant ElAgroudy, Yasmeen Abdrabou, Caroline Eckerth, Sarah Diefenbach, and Kristina Knaving. 2021. ‘I Don’t Need a Goal’: Attitudes and Practices in Fitness Tracking beyond WEIRD User Groups. In Proceedings of the 23rd International Conference on Mobile Human-Computer Interaction (Toulouse & Virtual, France) (MobileHCI ’21). Association for Computing Machinery, New York, NY, USA, Article 15, 14 pages. https://doi.org/10.1145/3447526.3472062
@@
- Orzikulova et al. (2024) Adiba Orzikulova, Han Xiao, Zhipeng Li, Yukang Yan, Yuntao Wang, Yuanchun Shi, Marzyeh Ghassemi, Sung-Ju Lee, Anind K Dey, and Xuhai Xu. 2024. Time2Stop: Adaptive and Explainable Human-AI Loop for Smartphone Overuse Intervention. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 250, 20 pages. https://doi.org/10.1145/3613904.3642747
@@
- Park and Lee (2023) Joonyoung Park and Uichin Lee. 2023. Understanding Disengagement in Just-in-Time Mobile Health Interventions. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 2, Article 72 (June 2023), 27 pages. https://doi.org/10.1145/3596240
@@
- Pina et al. (2017) Laura R. Pina, Sang-Wha Sien, Teresa Ward, Jason C. Yip, Sean A. Munson, James Fogarty, and Julie A. Kientz. 2017. From Personal Informatics to Family Informatics: Understanding Family Practices around Health Monitoring. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing (Portland, Oregon, USA) (CSCW ’17). Association for Computing Machinery, New York, NY, USA, 2300–2315. https://doi.org/10.1145/2998181.2998362
@@
- Raj et al. (2019) Shriti Raj, Joyce M. Lee, Ashley Garrity, and Mark W. Newman. 2019. Clinical Data in Context: Towards Sensemaking Tools for Interpreting Personal Health Data. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 3, 1, Article 22 (March 2019), 20 pages. https://doi.org/10.1145/3314409
@@
- Rey et al. (2023) Bradley Rey, Bongshin Lee, Eun Kyoung Choe, and Pourang Irani. 2023. Investigating In-Situ Personal Health Data Queries on Smartwatches. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 6, 4, Article 179 (Jan. 2023), 19 pages. https://doi.org/10.1145/3569481
- Robinson et al. (2015) Charlotte Robinson, Clara Mancini, Janet van der Linden, Claire Guest, and Lydia Swanson. 2015. Exploring assistive technology for assistance dog owners in emergency situations. In Proceedings of the 8th ACM International Conference on PErvasive Technologies Related to Assistive Environments (Corfu, Greece) (PETRA ’15). Association for Computing Machinery, New York, NY, USA, Article 92, 2 pages. https://doi.org/10.1145/2769493.2769576
@@
- Shen et al. (2025) Vivian Shen, Xiaoying Yang, Chris Harrison, and Yang Zhang. 2025. Hapt-Aids: Self-Powered, On-Body Haptics for Activity Monitoring. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 129 (Sept. 2025), 26 pages. https://doi.org/10.1145/3749468
@@
- Sullivan and Lachman (2017) Alycia N Sullivan and Margie E Lachman. 2017. Behavior change with fitness technology in sedentary adults: a review of the evidence for increasing physical activity. Frontiers in public health 4, Article 289 (2017), 16 pages.
@@
- Wang et al. (2024) Jianyang Wang, Dongheng Zhang, Binbin Zhang, Jinbo Chen, Yang Hu, and Yan Chen. 2024. RF-GymCare: Introducing Respiratory Prior for RF Sensing in Gym Environments. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 3, Article 135 (Sept. 2024), 28 pages. https://doi.org/10.1145/3678568
@@
- Wolf (2019) Christine T. Wolf. 2019. Explainability scenarios: towards scenario-based XAI design. In Proceedings of the 24th International Conference on Intelligent User Interfaces (Marina del Ray, California) (IUI ’19). Association for Computing Machinery, New York, NY, USA, 252–257. https://doi.org/10.1145/3301275.3302317
@@
- Yang et al. (2015) Rayoung Yang, Eunice Shin, Mark W. Newman, and Mark S. Ackerman. 2015. When fitness trackers don’t ’fit’: end-user difficulties in the assessment of personal tracking device accuracy. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Osaka, Japan) (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 623–634. https://doi.org/10.1145/2750858.2804269
- Yfantidou et al. (2023) Sofia Yfantidou, Pavlos Sermpezis, Athena Vakali, and Ricardo Baeza-Yates. 2023. Uncovering Bias in Personal Informatics. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 3, Article 139 (Sept. 2023), 30 pages. https://doi.org/10.1145/3610914
@@
Appendix A Thematic Analysis Coding Chart
| ††: Thematic Analysis table. | ||
| Themes | Sub-Theme/Code | Design implication |
| Barriers to Proactivity | Ecosystem Fragmentation & Inter-op Problems Definition: Friction from multiple non-interoperable apps/brands and scattered records preventing longitudinal PHI use. Evidence (source): Interviews. | Design interoperable dashboards, import/export (EHR) connectors, and a single-view aggregator; prioritize Android-first flows based on market share. |
| Low Digital Health Literacy Definition: Users cannot interpret metrics or know what action to take from raw analytics. Evidence (source): Field notes and Interviews. | Provide literacy-appropriate threshold-based plain-language recommendations and visual metaphors. | |
| Trust & Sensor-Confidence Issues Definition: Doubts about the accuracy of advanced metrics; users cross-check multiple sources before acting. Evidence (source): Interviews and Co-design workshops | Add sensor-confidence indicators, provenance badges, cross-device comparison views, and clinician-validated thresholds. | |
| Device & Form-factor Friction Definition: Physical discomfort, charging burden, and undesired form factors cause abandonment. Evidence (source): Interviews | Support multiple modalities (bands, rings, app-only), low-maintenance devices, battery-friendly modes, and non-wearable capture options. | |
| Preference for Embodied/Traditional Cues & Reactive Norms Definition: Reliance on body-feeling and doctor visits; belief that reactive care is sufficient. Evidence (source): Interviews and Co-design workshops. | Design a hybrid onboarding that maps PHI outputs to familiar embodied cues; provide low-friction “check only when alerted” modes. | |
| Enablers of Proactivity: Behaviours & Triggers | Stakeholder specific sharing Definition: PHI used as a shared family resource; younger family members’ recommendations create derived trust; Linking PHI to clinicians enables earlier action. Evidence (source): Interviews and health-sharing survey: 114 respondents — 20 users, 54 non-users, 40 unaware. | Create a user-friendly family-sharing UI with approval/request features. Include one-page PDF/HL7-style exports, customizable clinician summaries, and time-based trend highlights. |
| Motivational Social Features Definition: Family challenges, badges, and friendly competition that encourage preventive routines. Evidence (source): Co-design workshops. | Add consensual family challenges, non-punitive badges, group goals, and shareable progress snapshots that respect privacy controls. | |
| Actionable Explanations (literacy-appropriate) Definition: Plain-language recommendations tied to metric thresholds (what to do, not just numbers). Evidence (source): Interviews and Co-design workshops. | Convert numbers actions: “If X over Y days, do Z”; contextual links to vetted local resources; in-app “explain this metric” with examples. | |
| Proactive Monitoring (behavior) Definition: Regular, routine checking of PHI and prevention-oriented actions (daily/weekly review, goal setting, early consultations). Evidence (source): 55/87 (60%) survey participants reported tracking with a wearable; interviews | Quantify proactive vs reactive segments; surface summarized trends (daily/weekly) and goal dashboards; surface family-monitoring affordances. | |
| Triggers for Proactivity Definition: Specific events, alerts, family prompts or clinician requests that cause earlier action (alerts, family messages, prototype “ask to share”). Evidence (source): Co-design workshops | Implement configurable alerts, “ask-to-share” nudges, permissioned push messages, and simple export/share buttons for clinician workflows. |